40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 19 08:37:21 2017, Jamie, Update, CDS, Update on front-end/DAQ rebuild  
    Reply  Wed Jul 19 14:26:50 2017, Jamie, Update, CDS, Update on front-end/DAQ rebuild  
    Reply  Fri Jul 21 18:03:17 2017, Jamie, Update, CDS, Update on front-end/DAQ rebuild  
       Reply  Sun Jul 23 22:16:55 2017, Jamie, gautam, Update, CDS, front-end now running with new OS, RCG 2017-07-23-210810_1394x488_scrot.png2017-07-23-211812_387x488_scrot.png
          Reply  Mon Jul 24 10:45:23 2017, gautam, Update, CDS, c1iscex models died c1iscexFailure.png
             Reply  Mon Jul 24 10:59:08 2017, Jamie, Update, CDS, c1iscex models died 
          Reply  Mon Jul 24 19:28:55 2017, Jamie, Update, CDS, front end MX stream network working, glitches in c1ioo fixed 48.png
             Reply  Mon Jul 24 19:57:54 2017, gautam, Update, CDS, IMC locked, Autolocker re-enabled 
             Reply  Wed Jul 26 19:13:07 2017, Jamie, Update, CDS, daqd showing same instability as before 
                Reply  Fri Jul 28 20:22:41 2017, Jamie, Update, CDS, possible stable daqd configuration with separate DC and FW 
                   Reply  Mon Jul 31 15:13:24 2017, gautam, Update, CDS, FB ---> FB1 
                   Reply  Mon Jul 31 18:44:40 2017, Jamie, Update, CDS, CDS system essentially fully recovered 02.png
                      Reply  Thu Aug 3 19:46:27 2017, Jamie, Update, CDS, new daqd restart procedure 
                      Reply  Fri Aug 4 09:07:28 2017, rana, Update, CDS, CDS system essentially NOT fully recovered 
                         Reply  Thu Aug 10 14:25:52 2017, gautam, Update, CDS, Slow EPICS channels -> Frames re-enabled 
                            Reply  Fri Aug 11 00:10:03 2017, gautam, Update, CDS, Slow EPICS channels -> Frames re-enabled 
                               Reply  Fri Aug 11 11:14:24 2017, gautam, Update, CDS, Slow EPICS channels -> Frames re-enabled 
                               Reply  Fri Aug 11 18:53:35 2017, gautam, Update, CDS, Slow EPICS channels -> Frames re-enabled 
                      Reply  Fri Aug 11 19:34:49 2017, Jamie, Update, CDS, CDS final bits status update 
Message ID: 13139     Entry time: Mon Jul 24 19:57:54 2017     In reply to: 13138
Author: gautam 
Type: Update 
Category: CDS 
Subject: IMC locked, Autolocker re-enabled 

Now that all the front end models are running, I re-aligned the IMC, locked it manually, and then tweaked the alignment some more. The IMC transmission now is hovering around 15300 counts. I re-enabled the Autolocker and FSS Slow loops on Megatron as well.

Quote:

MX/OpenMX network running

Today I got the mx/open-mx networking working for the front ends.  This required some tweaking to the network interface configuration for the diskless front ends, and recompiling mx and open-mx for the newer kernel.  Again, this will all be documented.

 

ELOG V3.1.3-